how many yards of material from a 24 yard length of cloth remains after 3 pieces each 31/2 yards long and 5 pieces each 2 1/4 long are removed/
Answer by liz2298(4)

Conversion Factor
1 yard = 3 feet
Length of Pieces Removed
3 1/2 yards = 3 1/2 yards (3 feet/1 yard) * 3 = 31 1/2 feet
2 1/4 yards = 2 1/4 yards (3 feet/1 yard) * 5 = 33 3/4 feet
31 1/2 + 33 3/4 = 65 1/4 feet
Convert Removed Length to Feet
65 1/4 feet (1 yard/3 feet) = 21 3/4 yards
Subtract Removed Length from Whole
24 yards – 21 3/4 yards = 23 4/4 yards – 21 3/4 yards = 2 1/4 yards